# AdaMuon: Adaptive Muon Optimizer

### v2

@@ 1 Introduction
Optimization algorithms are a cornerstone of modern deep learning, directly shaping training dynamics and influencing both convergence speed and generalization performance. As model scales have grown to billions or even trillions of parameters (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Liu et al., 2024; Moonshot, 2025), optimizers face increasingly heterogeneous gradient landscapes and complex parameter geometries. Adaptive methods (Duchi et al., 2011; Tieleman, 2012; Loshchilov & Hutter, 2017), exemplified by Adam (Kingma & Ba, 2014), have become the de facto choice for large-scale training owing to their variance-based step size control and ease of tuning. However, over a decade after its introduction, the field is witnessing a growing demand for a new generation of optimizers that are better aligned with the computational and statistical challenges of training modern large foundation models.
@@
In the remainder of this paper, Sec. 2 reviews Muon and its extensions, along with the necessity and challenges of incorporating second momentum estimation; Sec. 3 presents the AdaMuon algorithm in detail; and Sec. 4 demonstrates that AdaMuon consistently outperforms other methods across a variety of models and datasets. Overall, we show that AdaMuon is a versatile algorithm that scales effectively to large-scale models.
@@
 Initialize first-momentum $\mathbf{M}_{0}\leftarrow\mathbf{0}$, second-momentum $\mathbf{V}_{0}\leftarrow\mathbf{0}$
@@ 2.3.1 Necessity
However, this global orthogonalization does not model the local variance structure of individual parameters. In real-world training, gradient distributions are often highly anisotropic at the element level—some entries exhibit large fluctuations, while others remain consistently stable. Applying a uniform update magnitude in such cases risks overshooting in noisy coordinates and under-updating informative but low-variance ones. This mismatch might slow convergence, reduce stability, and limit Muon‘s effectiveness in tasks with heterogeneous gradient statistics.
@@ 3.1 Element-Wise Second Momentum Estimation
$\hat{\mathbf{O}}_{t}={\mathbf{O}_{t}}\oslash({\sqrt{\mathbf{V}_{t}}+\varepsilon\cdot\mathbf{1}}),$ (6)
@@ 3.2 Stabilizing Orthogonal Updates via Sign Transformation
We hope $\mathbf{O}_{t}$ to be both usable in second momentum scaling and stable enough for variance-based normalization. To achieve this, we propose to first apply a transformation $f$ to $\mathbf{M}_{t}$, so that
@@ 3.3 RMS-Aligned Rescaling
To maintain compatibility with Adam’s learning rate schedules, we scale the RMS norm of AdaMuon’s update $\hat{\mathbf{O}}_{t}$ (after second momentum estimation in Eq. (6)) to match Adam’s empirical RMS value of $\approx 0.2$ (Liu et al., 2025), yielding
$\gamma_{t}=\frac{0.2}{{\rm RMS}(\hat{\mathbf{O}}_{t})}=\frac{0.2\sqrt{mn}}{\|\hat{\mathbf{O}}_{t}\|_{F}}.$ (8)
@@
$\mathbf{W}_{t+1}=\mathbf{W}_{t}-\eta\big{(}\gamma_{t}\hat{\mathbf{O}}_{t}+\lambda\mathbf{W}_{t}\big{)},$ (9)
The pseudo-code of AdaMuon is shown in Alg. 1. In addition, we provide further discussions on the omission of bias correction in the second momentum estimation (unlike Adam). Please refer to Appendix B for more details.
@@ 4.1 Experimental Setup
We compare AdaMuon against AdamW (Loshchilov & Hutter, 2017) and Muon (Jordan et al., 2024). Given Muon’s strong performance in large-scale training, we consider these two baselines sufficient to demonstrate the effectiveness of AdaMuon. Since Muon operates only on matrices and applies a separate Adam optimizer for the remaining parameters, we set the learning rates of both optimizers to be the same. For AdamW, the first and second momentum coefficients are set to 0.9 and 0.95, respectively; Muon uses $\beta=0.95$, and AdaMuon introduces no extra hyper-parameters. The weight decay $\lambda$ for these optimizers are set to 0.1.
@@
We evaluate AdaMuon on two representative model families: GPT-2 and Qwen2.5. All GPT-2 experiments are based on the nanoGPT (Karpathy, 2022) implementation of the GPT-2 architecture (Radford et al., 2019). We consider four scales—Small (125M), Medium (355M), Large (770M), and XL (1.5B). Following the default configurations in nanoGPT, we remove bias terms in all linear layers, use the GeLU activation function, and set the dropout rate to 0.0. Two modifications are applied: (1) replacing the original learned positional embedding (WPE) with Rotary Positional Embedding (RoPE) (Su et al., 2024), and (2) substituting the cosine learning rate schedule with the warmup-stable policy. For Qwen2.5 (Qwen et al., 2025), we adopt the 1.5B dense model following the official architecture specifications.
@@
GPT-2 models are trained on the OpenWebText dataset (Gokaslan et al., 2019), which contains approximately 9B training tokens and 4.4M validation tokens, all tokenized with the standard GPT-2 tokenizer. For Qwen2.5 training, the dataset is collected from online corpora, with low-quality content removed through a combination of mannually filtering rules and LLM-based quality assessment. Beyond general text, it contains high-quality code, math, and multilingual content.
@@
For GPT-2 experiments, all models are trained on approximately 49.2B training tokens for 100K steps, with a context length of 1024 and a warmup period of 2K steps. For Qwen2.5 experiments, the model is trained on 100B tokens with a context length of 8196 with 6e-4 learning rate. All the other hyper-parameters follow the respective official configurations. All other training settings including random seed, etc, unless otherwise specified, are kept consistent across methods.
Evaluation Metric
4.2 Result
GPT-2 Results.
Fig. 1 illustrates the loss–token curves of AdamW, Muon and AdaMuon under two learning-rate settings, and Table 1 summarizes the corresponding training efficiency of Muon and AdaMuon. Across all four GPT-2 scales, both Muon and AdaMuon deliver significant efficiency gains over AdamW, validating the benefit of geometry-preserving updates. Notably, AdaMuon consistently achieves the highest efficiency, regardless of model size or learning rate, demonstrating the effectiveness of integrating second-moment scaling with RMS-norm alignment. We also note that for GPT-2 XL with a learning rate of $1\times 10^{-3}$, the efficiency gap is particularly large: in this regime, AdamW exhibits unstable training with multiple loss spikes, whereas Muon and AdaMuon remain stable, enabling them to reach the target loss substantially faster.
@@ 4.2 Result
Figure 2: Results of AdamW, Muon, and AdaMuon when training Qwen2.5-1.5B dense model.
Table 2: Evaluation of Qwen2.5-1.5B trained by AdamW, Muon and AdaMuon. Abbreviations: ARC-C = ARC-Challenge, WinoG = WinoGrande, ComQA = CommonsenseQA.
Optimizer MMLU MMLU-Pro TriviQA ARC-C GPQA OBQA HellaSwag WinoG
@@
Optimizer PIQA ComQA CHID CMMLU CEval MBPP GSM8k Avg.
@@
Qwen2.5 Results.
@@ Appendix A Proof of Theorem 1
Appendix B Rationale for Omitting Bias Correction in AdaMuon’s Second Momentum
@@
$\hat{\mathbf{V}}_{t}=\frac{\mathbf{V}_{t}}{1-\beta^{t}}$ (11)
@@
$\hat{\mathbf{O}}_{t}=\frac{\mathbf{O}_{t}}{\sqrt{\mathbf{V}_{t}}+\varepsilon}$ (12)
@@
$\tilde{\mathbf{O}}_{t}=\gamma\,\hat{\mathbf{O}}_{t},\quad\gamma=\frac{0.2\cdot\sqrt{mn}}{\|\hat{\mathbf{O}}_{t}\|_{F}},$ (13)
@@
$\hat{\mathbf{O}}_{t}\propto\frac{1}{\sqrt{c}},\quad\gamma\propto\sqrt{c},$ (14)

### v3

@@ 1 Introduction
Optimization algorithms are a cornerstone of modern deep learning, directly shaping training dynamics and influencing both convergence speed and generalization performance. As model scales have grown to billions or even trillions of parameters (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Liu et al., 2024b; Moonshot, 2025), optimizers face increasingly heterogeneous gradient landscapes and complex parameter geometries. Adaptive methods (Duchi et al., 2011; Tieleman, 2012; Loshchilov & Hutter, 2017), exemplified by Adam (Kingma & Ba, 2014), have become the de facto choice for large-scale training owing to their variance-based step size control and ease of tuning. However, over a decade after its introduction, the field is witnessing a growing demand for a new generation of optimizers that are better aligned with the computational and statistical challenges of training modern large foundation models.
@@
In the remaining of this paper, Sec. 2 reviews Muon and its extensions, along with the necessity and challenges of incorporating second momentum estimation; Sec. 3 presents the AdaMuon algorithm in detail; and Sec. 4 demonstrates that AdaMuon consistently outperforms other methods across a variety of models and datasets. Overall, we show that AdaMuon is a versatile algorithm that scales effectively to large-scale models.
@@
 Initialize first momentum $\mathbf{M}_{0}\leftarrow\mathbf{0}$, second momentum $\mathbf{V}_{0}\leftarrow\mathbf{0}$
@@ 2.3.1 Necessity
However, this global orthogonalization does not model the local variance structure of individual parameters. In real-world training, gradient distributions are often highly anisotropic at the element level—some entries exhibit large fluctuations, while others remain consistently stable. Applying a uniform update magnitude in such cases risks overshooting in noisy coordinates and under-updating informative but low-variance ones. This mismatch might slow convergence, reduce stability, and limit Muon’s effectiveness in tasks with heterogeneous gradient statistics.
@@ 3.1 Element-Wise Second Momentum Estimation
$\widehat{\mathbf{O}}_{t}={\mathbf{O}_{t}}\oslash({\sqrt{\mathbf{V}_{t}}+\varepsilon\cdot\mathbf{1}}),$ (6)
@@ 3.2 Stabilizing Orthogonal Updates via Sign Transformation
We hope that $\mathbf{O}_{t}$ can be both utilized in second momentum scaling and stable enough for variance-based normalization. To achieve this, we propose to first apply a transformation $f$ to $\mathbf{M}_{t}$, so that
@@ 3.3 RMS-Aligned Rescaling
To maintain compatibility with Adam’s learning rate schedules, we scale the RMS norm of AdaMuon’s update $\widehat{\mathbf{O}}_{t}$ (after second momentum estimation in Eq. (6)) to match Adam’s empirical RMS value of $\approx 0.2$ (Liu et al., 2025), yielding
$\gamma_{t}=\frac{0.2}{{\rm RMS}(\widehat{\mathbf{O}}_{t})}=\frac{0.2\sqrt{mn}}{\|\widehat{\mathbf{O}}_{t}\|_{F}}.$ (8)
@@
$\mathbf{W}_{t+1}=\mathbf{W}_{t}-\eta\big(\gamma_{t}\widehat{\mathbf{O}}_{t}+\lambda\mathbf{W}_{t}\big),$ (9)
The pseudo-code of AdaMuon is shown in Alg. 1. In addition, we provide further discussions on the omission of bias correction in the second momentum estimation (Appendix B), and the convergence analysis of AdaMuon (Appendix C).
@@ 4.1 Experimental Setup
We compare AdaMuon against AdamW (Loshchilov & Hutter, 2017) and Muon (Jordan et al., 2024). Given Muon’s strong performance in large-scale training, we consider these two baselines sufficient to demonstrate the effectiveness of AdaMuon. Muon here is specifically the Kimi-variant Liu et al. (2025). Since Muon operates only on matrices and applies a separate Adam optimizer for the remaining parameters, we set the learning rates of both optimizers to be the same. For AdamW, the first and second momentum coefficients are set to 0.9 and 0.95, respectively; Muon uses $\beta=0.95$, and AdaMuon uses $\beta=0.95$ and $\epsilon=10^{-8}$. The weight decay $\lambda$ for these optimizers are set to 0.1.
@@
We evaluate AdaMuon on two representative model families: GPT-2 and Qwen2.5 models. All GPT-2 experiments are based on the nanoGPT (Karpathy, 2022) implementation of the GPT-2 architecture (Radford et al., 2019). We consider four scales—Small (125M), Medium (355M), Large (770M), and XL (1.5B). Following the default configurations in nanoGPT, we remove bias terms in all linear layers, use the GeLU activation function, and set the dropout rate to 0.0. Two modifications are applied: (1) replacing the original learned positional embedding (WPE) with Rotary Positional Embedding (RoPE) (Su et al., 2024), and (2) substituting the cosine learning rate schedule with the warmup-stable policy (i.e., the schdule that omits the decay phase in WSD (Hu et al., 2024) schedule entirely: after warmup, the learning rate is held constant). For Qwen2.5 (Qwen et al., 2025), we adopt the 1.5B and 7B dense model following the official architecture specifications.
@@
GPT-2 models are trained on the OpenWebText dataset (Gokaslan et al., 2019), which contains approximately 9B training tokens and 4.4M validation tokens, all tokenized with the standard GPT-2 tokenizer. For Qwen2.5 models, the dataset is collected from online corpora, with low-quality content removed through a combination of manually filtering rules and LLM-based quality assessment. Beyond general text, it contains high-quality code, math, and multilingual content.
@@
We mainly focus on the pre-training of the model. For GPT-2 experiments, all models are trained on approximately 50B (49.2B) training tokens for 100K steps, with a context length of 1024 and a warmup period of 2K steps. For Qwen2.5 experiments, 1.5B model is trained on 100B tokens with a context length of 8196 with 6e-4 learning rate, and 7B model is trained on 235B tokens. For 1.5B model, we use 8.4B tokens for warmup, and 16.8B for 7B model. All the other hyper-parameters follow the respective official configurations. All other training settings, unless otherwise specified, are kept consistent across methods.
For GPT runs, we fix the effective batch size (EBS) to 60 sequences. By model size (Small to XL), we use the following pairs (batch size (bs), gradient accumulate steps (gas)) so that $\text{EBS}=\text{bs}\times\text{gas}$ = 60: (15, 4), (15, 4), (5, 12), (5, 12). For Qwen-1.5B, we use global batch size (gbs) = 512 with micro-batch size (mbs) = 2; for Qwen-7B, gbs = 1024, mbs = 2.
@@
Evaluation Metric.
Evaluation Benchmark.
We evaluate the Qwen model on a broad spectrum of 15 benchmarks. For English language understanding and reasoning, we adopt MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), TriviQA (Joshi et al., 2017), ARC-C (Clark et al., 2018), GPQA (Rein et al., 2024), OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), and CommomsenseQA (Talmor et al., 2018). For code generation, we evaluate on the MBPP (Austin et al., 2021). For mathematical reasoning, we use the GSM8k (Cobbe et al., 2021). For Chinese language understanding and reasoning, we include CHID (Zheng et al., 2019), CMMLU (Li et al., 2023), and CEval (Huang et al., 2023).
4.2 Result
GPT-2 Results.
Table 1 summarizes the training efficiency of Muon and AdaMuon under two learning-rate settings, and Fig. 1 illustrates the corresponding loss–token curves of AdamW, Muon and AdaMuon. Across all four GPT-2 scales, both Muon and AdaMuon deliver significant efficiency gains over AdamW, validating the benefit of geometry-preserving updates. Notably, AdaMuon consistently achieves the highest efficiency, regardless of model size or learning rate, demonstrating the effectiveness of integrating second-moment scaling with RMS-norm alignment. We also note that for GPT-2 XL with a learning rate of $1\times 10^{-3}$, the efficiency gap is particularly large: in this regime, AdamW exhibits unstable training with multiple loss spikes, whereas Muon and AdaMuon remain stable, enabling them to reach the target loss substantially faster
Figure 2: Results of AdamW, Muon, and AdaMuon when training Qwen2.5-1.5B and 7B dense models.
Qwen2.5 Results.
Table 2: Evaluation of Qwen2.5 models trained by AdamW, Muon and AdaMuon. Abbreviations: ARC-C = ARC-Challenge, WinoG = WinoGrande, ComQA = CommonsenseQA.
Model Optimizer MMLU MMLU-Pro TriviaQA ARC-C GPQA OBQA HellaSwag WinoG
@@ 4.2 Result
7B AdamW 33.30 7.44 26.16 33.33 28.00 44.44 44.59 58.07
Muon 34.63 6.73 30.42 24.32 28.00 56.35 43.31 58.49
AdaMuon 34.98 10.87 30.22 42.32 32.00 53.70 46.50 58.20
Model Optimizer PIQA ComQA CHID CMMLU CEval MBPP GSM8k Avg.
@@
7B AdamW 73.91 32.68 44.32 31.12 32.49 8.80 5.46 33.61
Muon 76.52 34.15 60.91 32.34 35.05 11.20 10.16 36.17
AdaMuon 72.90 29.25 58.26 32.31 35.86 13.20 10.08 37.38
4.3 Ablation Study
Figure 3: Ablation Study of AdaMuon. We present the training loss curve of GPT-2 Small and Qwen2.5-1.5B models.
We conduct ablation experiments by selectively removing the sign operation and the second momentum term, with results summarized in Fig. 3. It can be observed that retaining only the sign operation still yields lower loss than the original Muon optimizer. This improvement stems from the fact that sign enhances the stability of the matrix obtained through polar decomposition, thereby stabilizing the update direction. In contrast, when only the second momentum term is added, the advantage disappears and Muon even outperforms this variant. This phenomenon is intuitive: directly accumulating second momentum can partly provide normalization, but since the orthogonal updates themselves remain unstable, the accumulated variance fails to stabilize training and may even harm optimization, leading to inferior loss reduction. Finally, when both sign and second momentum are retained, they complement each other: the former improves stability of the orthogonal update, while the latter provides effective element-wise normalization, resulting in the best overall performance.
4.4 Training Behavior
In this subsection, we further analyze the training behavior of the three optimizers. All experiments are conducted on GPT-2 1.5B with a learning rate of $6\times 10^{-4}$. We examine their gradient norms, parameter update norms, and the evolution of max attention logits across layers.
As shown in Fig. 4, the gradient norms of the three optimizers follow different trajectories. Muon exhibits a gradual upward drift, reflecting increasing gradient variance as training progresses. AdamW stabilizes quickly, maintaining a relatively flat trajectory with mild fluctuations, which indicates effective variance control. AdaMuon behaves similarly to AdamW but produces even smoother updates, highlighting its stronger capability in suppressing short-term oscillations.
The parameter norms further highlight their divergent behaviors. Both Muon and AdamW eventually converge to similar magnitudes, with Muon showing a sharper initial rise, consistent with its more aggressive gradient scaling. AdaMuon consistently produces smaller parameter norm, converging to a noticeably lower parameter norm. This restrained behavior helps stabilize training, although it may reduce the optimizer’s ability to fully exploit the model’s capacity compared to Muon or Adam.
Finally, the per-layer max attention logits reveal distinct adaptation patterns. In shallow layers ( Layer 4), Muon steadily reduces the maximum attention logits, while AdamW maintains relatively high values, indicating stronger localized activations. In mid layers (Layer 25), all optimizers reduce logits, but AdaMuon stabilizes earlier, whereas Muon continues a slow decline. In deeper layers (Layer 45), AdamW sustains substantially higher logits than the others, while both Muon and AdaMuon settle at lower levels. These results suggest that AdamW encourages deeper specialization of attention heads, whereas Muon and AdaMuon promote a more uniform redistribution of attention across layers.
Figure 4: Training behavior of AdamW, Muon, and AdaMuon.
In addition, we note that prior work (Liu et al., 2025) reported that Muon tends to increase the maximum attention logits. Our findings differ: in the standard multi-head attention (MHA) setting, we do not observe excessively large logits under Muon. We argue that this discrepancy is not caused by the optimizer itself, but rather by the model architecture. Specifically, those studies employed variants with multi-latent attention (Liu et al., 2024a), where the structural design inherently amplifies logit magnitudes. By contrast, in conventional MHA layers, the effect of Muon on maximum logits remains moderate and comparable to other optimizers.
4.5 Sensitivity Analysis
Table 3: Final training loss of varied $\beta$ on GPT2-small.
$\beta$ 0.8 0.9 0.95 0.99
Loss 2.96107 2.95735 2.95521 2.95487
We evaluate AdaMuon’s sensitivity to the first- and second-momentum coefficient $\beta$ on GPT-small with learning rate $6\times 10^{-4}$. Varying $\beta$ over a standard range, the final-step training losses are shown in Table 3, which cluster tightly, indicating low sensitivity as long as $\beta$ remains in a reasonable band. Therefore, no special tuning is required—AdaMuon can directly reuse Muon’s default $\beta$ (0.95).
@@ 5 Conclusion
6 Limitation
AdaMuon also has some limitations. First, one additional second-moment buffer introduces more computation and memory. Second, there is minor per-step runtime overhead from Newton–Schulz and variance EMA. How to effectively tackle these drawbacks is also interesting, leaving for the future work.
@@ Appendix A Proof of Theorem 1
Appendix B Why Omitting Bias Correction in AdaMuon’s Second Momentum
@@
$\widehat{\mathbf{V}}_{t}=\frac{\mathbf{V}_{t}}{1-\beta^{t}}$ (11)
@@
$\widehat{\mathbf{O}}_{t}={\mathbf{O}_{t}}\oslash({\mathbf{V}_{t}+\varepsilon\mathbf{1}})$ (12)
@@
$\tilde{\mathbf{O}}_{t}=\gamma\,\widehat{\mathbf{O}}_{t},\quad\gamma=\frac{0.2\cdot\sqrt{mn}}{\|\widehat{\mathbf{O}}_{t}\|_{F}},$ (13)
@@
$\widehat{\mathbf{O}}_{t}\propto\frac{1}{\sqrt{c}},\quad\gamma\propto\sqrt{c},$ (14)